# From Automation to Autonomy: A Survey on Large Language Models in Scientific Discovery

### v1

@@ 1 Introduction
However, this rapid expansion of LLM applications and the ongoing paradigm shift in scientific discovery present notable challenges. The accelerated pace of LLM evolution and their deepening integration into complex research complicate systematic assessment, necessitating conceptual frameworks to structure current understanding and chart future directions. While existing surveys have provided valuable overviews of LLMs in various scientific domains Zhang et al. (2024, 2025) or have cataloged particular AI techniques for science Luo et al. (2025); Reddy and Shojaee (2025), they often focus on discipline-specific applications or a static snapshot of LLM capabilities. Consequently, existing reviews may overlook the crucial trend of increasing LLM autonomy and their evolving roles across the entire scientific method, leaving their comprehensive impact and trajectory towards greater independence underexplored.
@@
This survey focuses on LLM-based systems in scientific discovery, particularly their varying levels of autonomy. While acknowledging the broad impact of LLMs in science, we deliberately narrow our scope to exclude research on general-purpose scientific LLMs or LLMs for domain-specific scientific knowledge acquisition and reasoning, which are well covered in existing surveys Zhang et al. (2024, 2025). The remainder of this paper is organized as follows: Section 2 details our taxonomy and its interaction with the scientific method. Section 3 presents LLM as Tool applications, categorized by scientific method stages. Section 4 examines LLM as Analyst works by scientific domain, while Section 5 analyzes LLM as Scientist systems, focusing on their idea development and refinement strategies. Section 6 explores challenges and future directions.
@@ 2 Three Levels of Autonomy
Collectively, we present our full taxonomy with detailed categorization in Figure 2, which consolidates 90 research works within our focused scope across all three levels of autonomy.
@@ 3.1 Literature Review and Information Gathering
Automatic literature search and retrieval is crucial for identifying research gaps and formulating research questions. Lála et al. (2023) first introduced the literature retrieval benchmark LitQA, featuring a RAG-based agent, PaperQA. LitLLM Agarwal et al. (2024) further provided a comprehensive RAG-based toolkit for LLM-driven literature review. Dennstädt et al. (2024) directed their focus to the biomedical domain, highlighting the potential of LLMs in literature screening. Other approaches, such as SCIMON Wang et al. (2024a) and ResearchAgent Baek et al. (2025), have integrated active literature retrieval with the generation of research ideas. More recently, Gao et al. (2025) tackled the task of hierarchically organizing scientific literature through fine-grained paper abstraction. Nevertheless, several ‘Deep Research’ products OpenAI (2025); Google (2025); xAI (2025) have recently been released, featuring enhanced agentic workflows that support automated literature web search, organization, and report generation, thereby significantly accelerating traditional, human-intensive literature research processes.
@@
In parallel, several works have explored effective methods for aggregating information from scientific papers into tabular summaries. ArxivDIGESTables Newman et al. (2024) investigated cross-literature table generation using LLMs, accompanied by an automated evaluation strategy. ArXiv2Table Wang et al. (2025) revised the evaluation protocol and provided a comprehensive benchmark. Methodologies such as Text-Tuple-Table Deng et al. (2024b) and TKGT Jiang et al. (2024b) have enhanced the quality of LLM-based table generation by incorporating tuple-based structures and graph modalities.
@@ 3.2 Idea Generation and Hypothesis Formulation
Numerous research efforts have focused on the automated generation of novel research ideas and hypotheses. In the general domain, benchmarks such as IdeaBench Guo et al. (2024a) and LiveIdeaBench Ruan et al. (2025) have evaluated the capability of LLMs to generate research ideas based on provided literature summaries. Concurrently, LLM-based agent frameworks, including Nova Hu et al. (2024a), SciAgents Ghafarollahi and Buehler (2024b), and KG-CoI Xiong et al. (2024), have been proposed to enhance idea generation through effective reasoning over academic knowledge graphs, iterative planning, and searching. Moreover, several exploratory studies have assessed the novelty and quality of LLM-generated ideas for AI research, underscoring the potential for automated idea generation when coupled with appropriate human guidance Si et al. (2024); Feng et al. (2025); Qiu et al. (2025). Furthermore, many studies within natural science disciplines have investigated LLM-based idea generation in domain-specific contexts. For example, Ciucă et al. (2023) proposed adopting adversarial prompting for effective idea generation in astronomy. In biology, Buehler (2024) enhanced idea generation by integrating knowledge extraction and graph representations.
@@
Building upon identified ideas, the design of testable scientific hypotheses has also been a significant focus. Qi et al. (2023) and Yang et al. (2024) examined the ability of LLMs to propose hypotheses, demonstrating their considerable capacity for generating novel yet valid hypotheses under open-ended constraints. Methodologies such as Scideator Radensky et al. (2025) have been developed to investigate human-LLM collaboration to facilitate grounded idea and hypothesis generation. Within natural science, benchmarks Qi et al. (2024) and methods O’Brien et al. (2024) have extended hypothesis generation into the biomedical domain. Meanwhile, MOOSE-Chem Yang et al. (2025) offers a systematic evaluation benchmark and an agent framework specifically for hypothesis discovery in chemistry.
@@ 3.4 Data Analysis and Organization
In this stage, LLMs assist the scientific workflow by automating processes related to data organization, presentation, and analysis. For data presented in tabular format, Chain-of-Table Wang et al. (2024c) proposes a method to enhance tabular understanding by incorporating evolving tables within the reasoning chain of LLMs. Concurrently, Deng et al. (2024a) highlight the potential of integrating visual information to improve multimodal understanding, thereby aiding tabular comprehension. More recently, Wu et al. (2025) introduced TableBench, a comprehensive benchmark for table-based question answering under practical industrial scenarios.
@@
Beyond tabular data, charts represent another important format for organizing and storing information derived from experimental data. Early benchmarks, exemplified by ChartQA Masry et al. (2022), examined the capabilities of vision transformers in chart-based question answering. Subsequent works, including CharXiv Wang et al. (2024d) and ChartX Xia et al. (2025), have expanded the scope of chart understanding scenarios by utilizing human-curated chart generation or by incorporating real-world chart data sourced from arXiv preprints. Regarding chart generation, AutomaTikZ Belouadi et al. (2024) formulates the process as TikZ code generation from caption text and has demonstrated the efficacy of fine-tuning LLMs using open scientific figure data. More recently, Text2Chart31 Zadeh et al. (2025) employed reinforcement learning with automated feedback to refine chart generation capabilities within the Matplotlib library.
@@ 3.5 Conclusion and Hypothesis Validation
In this context, a significant focus of contemporary research involves investigating the utility of LLMs as reviewers for artificial intelligence papers. ReviewerGPT Liu and Shah (2023) initially explored the capability of LLMs to identify deliberately inserted errors within research papers, highlighting the necessity for more robust systems to conduct comprehensive reviews. Zhou et al. (2024a) further evaluated static LLMs in the context of reviewing real-world conference papers using a multiple-choice format. Du et al. (2024) conducted a comprehensive analysis of LLM review quality through extensive human studies and comparisons, revealing weaknesses in their ability to identify deficiencies. ClaimCheck Ou et al. (2025) further investigated the capabilities of LLMs in critiquing research claims, demonstrating that this task remains challenging even for highly advanced models such as OpenAI’s o1 OpenAI (2024). Concurrently, research highlights the potential to address these limitations by incorporating multi-agent systems with specialized roles Tyser et al. (2024); Xu et al. (2025a) or through LLM alignment via reinforcement learning Weng et al. (2025).
@@
Another important application at this stage is the automatic validation of hypotheses by LLMs. Takagi et al. (2023) demonstrated that LLMs possess considerable capabilities in automatically generating code to verify research hypotheses within simplified machine learning problems. SciReplicate-Bench Xiang et al. (2025) further extended this concept to the replication of real-world research papers. Furthermore, Xu et al. (2025c) have navigated this domain into physics research, aiming to enhance the interpretability of the discovery process through the use of multi-agent workflows.
@@ 4.1 Machine Learning Research
Automated Machine Learning (AutoML) Shen et al. (2024) endeavors to generate high-performing modeling configurations for a given task in a data-driven manner. With the advent of LLM-based agents, several studies have explored their application in the automated modeling of machine learning (ML) tasks. MLAgentBench Huang et al. (2024a) evaluates the capabilities of LLMs in designing and executing ML experiments, revealing that LLM performance is often contingent upon task familiarity derived from their training data. The IMPROVE framework Xue et al. (2025) highlighted the significance of iterative refinement mechanisms within LLM agents for enhancing modeling outcomes. CodeScientist Jansen et al. (2025) incorporated an ML modeling agent with machine-generated ideas, thereby progressing towards more autonomous ML research aimed at performance enhancement. In a related development, BudgetMLAgent Gandhi et al. (2025) leverages curated expert collaboration frameworks to achieve superior results using considerably more cost-effective open-source models.
@@ 4.2 Data Modeling and Analysis
Automated data-driven analysis, encompassing statistical data modeling and hypothesis validation, represents a foundational application area for LLM-assisted scientific discovery. InfiAgent-DABench Hu et al. (2024b) benchmarked the capabilities of LLMs in static code generation and execution for data analysis using CSV files. Subsequent benchmarks, such as BLADE Gu et al. (2024) and DiscoveryBench Majumder et al. (2024), have improved evaluation robustness by incorporating real-world research papers and expert-curated analytics within their benchmarking environments. These studies indicate that most LLMs struggle with complex data analytics tasks, even when operating within an agent framework Zheng et al. (2023). To address these challenges, DS-Agent Guo et al. (2024b) proposes to enhance LLM performance by incorporating a case-based reasoning method to improve domain knowledge acquisition. In a related effort, DAgent Xu et al. (2025b) extended the application domain to querying relational databases and enabled report generation using results derived from decomposed sub-problems.
@@ 4.3 Function Discovery
Function discovery, which aims to identify the underlying equations from observational data of variables, has been significantly influenced by the advancement of AI-driven symbolic regression (SR) Udrescu and Tegmark (2020); Kamienny et al. (2022). LLM-SR Shojaee et al. (2025a) enhanced function discovery by leveraging the prior domain knowledge of LLMs and incorporating feedback from clustered memory storage. LLM-SRBench Shojaee et al. (2025b) introduced a benchmark for evaluating LLMs as function discovery agents, which incorporates function transformations to mitigate data contamination. Furthermore, other studies have explored the capabilities of LLMs in discovering complex models within specific domains, such as Physics Koblischke et al. (2025) and Statistics Li et al. (2024b).
@@ 4.4 Natural Science Research
Research has largely focused on applying LLMs to autonomous research workflows for natural science discovery. Auto-Bench Chen et al. (2025a) evaluated LLMs on chemistry and social science tasks based on causal graph discovery, revealing that LLMs perform effectively only when task complexity is highly limited. In contrast, ScienceAgentBench Chen et al. (2025b) provided a multidisciplinary benchmark for LLMs operating within agent frameworks such as CodeAct Wang et al. (2024b) and self-debug Chen et al. (2023). This benchmark highlighted the necessity for tailored agent workflows for such explorative tasks.
@@ 4.5 General Research
Apart from specialized domain applications, some benchmarks have broadly evaluated diverse tasks from different stages of scientific discovery. DiscoveryWorld Jansen et al. (2024) created a virtual environment for LLM agents to conduct simplified scientific exploration. In Liu et al. (2025a), various application scenarios for AI agents in research were comprehensively discussed, supported by preliminary evaluation datasets. CURIE Cui et al. (2025) provides a comprehensive benchmark covering tasks in long-context understanding, reasoning, and information extraction, all of which are essential for scientific research. While EAIRA Cappello et al. (2025) focused on assessing the ability of LLMs to perform in a real-world research assistant role using various task formats.
@@ 6 Challenges and Future Directions
As LLM-based systems gain independent reasoning and action capabilities, their potential for risks—ranging from amplified societal biases to deliberate misuse like generating harmful substances or challenging human control—becomes increasingly salient and complex He et al. (2023); Ahadian and Guan (2024); Bengio et al. (2025). With AI capabilities and societal norms in constant flux, alignment is consequently an imperative, continuous process demanding adaptive governance and evolving value systems Li et al. (2024c). This requires embedding ethical constraints directly in scientific AI design frameworks, alongside vigilant oversight, to ensure advancements serve human well-being and the common good.
@@
Consequently, the scope of this survey has certain limitations. Firstly, we do not provide an exhaustive review of research focused on the development or fine-tuning of general-purpose scientific LLMs primarily designed for domain-specific knowledge acquisition, reasoning, or application. These areas, while crucial to the broader landscape of AI in science, are extensively covered in other existing surveys and fall outside our specific focus on the autonomy paradigm. Secondly, while we acknowledge the importance of fundamental LLM capabilities such as planning, code generation, and agentic decision-making, this survey does not delve deeply into orthogonal benchmarks or methodologies related to these general abilities unless they are explicitly contextualized and applied within the framework of LLM-driven scientific discovery and its evolving autonomy. These exclusions were deliberate to maintain a focused exploration of the transformative roles and increasing independence of LLMs throughout the scientific research lifecycle, rather than to offer a comprehensive catalog of all LLM-related tools or foundational capabilities that might be beneficial to scientific endeavors.

### v3

@@ 1 Introduction
However, this rapid expansion of LLM applications and the ongoing paradigm shift in scientific discovery present notable challenges. The accelerated pace of LLM evolution and their deepening integration into complex research complicate systematic assessment, necessitating conceptual frameworks to structure current understanding and chart future directions. While existing surveys have provided valuable overviews of LLMs in various scientific domains Zhang et al. (2024, 2025a) or have cataloged particular AI techniques for science Luo et al. (2025); Reddy and Shojaee (2025), they often focus on discipline-specific applications or a static snapshot of LLM capabilities. Consequently, existing reviews may overlook the crucial trend of increasing LLM autonomy and their evolving roles across the entire scientific method, leaving their comprehensive impact and trajectory towards greater independence underexplored.
@@
This survey focuses on LLM-based systems in scientific discovery, particularly their varying levels of autonomy. While acknowledging the broad impact of LLMs in science, we deliberately narrow our scope to exclude research on general-purpose scientific LLMs or LLMs for domain-specific scientific knowledge acquisition and reasoning, which are well covered in existing surveys Zhang et al. (2024, 2025a). The remainder of this paper is organized as follows: Section 2 details our taxonomy and its interaction with the scientific method. Section 3 presents LLM as Tool applications, categorized by scientific method stages. Section 4 examines LLM as Analyst works by scientific domain, while Section 5 analyzes LLM as Scientist systems, focusing on their idea development and refinement strategies. Section 6 explores challenges and future directions.
@@ 2 Three Levels of Autonomy
Collectively, we present our full taxonomy with detailed categorization in Figure 2, which consolidates research works within our focused scope across all three levels of autonomy.
@@ 3.1 Literature Review and Information Gathering
Automatic literature search and retrieval is crucial for identifying research gaps and formulating research questions. Lála et al. (2023) first introduced the literature retrieval benchmark LitQA, featuring a RAG-based agent, PaperQA. LitLLM Agarwal et al. (2024) further provided a comprehensive RAG-based toolkit for LLM-driven literature review. Taking this automation a step further, Wang et al. (2024c) demonstrated that large language models can automatically write entire survey papers. Dennstädt et al. (2024) directed their focus to the biomedical domain, highlighting the potential of LLMs in literature screening. Other approaches, such as SCIMON Wang et al. (2024a) and ResearchAgent Baek et al. (2025), have integrated active literature retrieval with the generation of research ideas. More recently, Gao et al. (2025) tackled the task of hierarchically organizing scientific literature through fine-grained paper abstraction. Nevertheless, several ‘Deep Research’ products OpenAI (2025); Google (2025); xAI (2025) have recently been released, featuring enhanced agentic workflows that support automated literature web search, organization, and report generation, thereby significantly accelerating traditional, human-intensive literature research processes.
@@
In parallel, several works have explored effective methods for aggregating information from scientific papers into tabular summaries. ArxivDIGESTables Newman et al. (2024) investigated cross-literature table generation using LLMs, accompanied by an automated evaluation strategy. ArXiv2Table Wang et al. (2025b) revised the evaluation protocol and provided a comprehensive benchmark. Methodologies such as Text-Tuple-Table Deng et al. (2024b) and TKGT Jiang et al. (2024b) have enhanced the quality of LLM-based table generation by incorporating tuple-based structures and graph modalities.
@@ 3.2 Idea Generation and Hypothesis Formulation
Numerous research efforts have focused on the automated generation of novel research ideas and hypotheses. In the general domain, benchmarks such as IdeaBench Guo et al. (2024a) and LiveIdeaBench Ruan et al. (2025) have evaluated the capability of LLMs to generate research ideas based on provided literature summaries. Concurrently, LLM-based agent frameworks, including Nova Hu et al. (2024a), SciAgents Ghafarollahi and Buehler (2024b), and KG-CoI Xiong et al. (2024), have been proposed to enhance idea generation through effective reasoning over academic knowledge graphs, iterative planning, and searching. More specific methodologies, such as employing dynamic control to guide the creative process, have also been introduced Li et al. (2024c). Moreover, several exploratory studies have assessed the novelty and quality of LLM-generated ideas for AI research, underscoring the potential for automated idea generation when coupled with appropriate human guidance Si et al. (2024); Feng et al. (2025); Qiu et al. (2025). Furthermore, many studies within natural science disciplines have investigated LLM-based idea generation in domain-specific contexts. For example, Ciucă et al. (2023) proposed adopting adversarial prompting for effective idea generation in astronomy. In biology, Buehler (2024) enhanced idea generation by integrating knowledge extraction and graph representations.
@@
Building upon identified ideas, the design of testable scientific hypotheses has also been a significant focus. Qi et al. (2023) and Yang et al. (2024) examined the ability of LLMs to propose hypotheses, demonstrating their considerable capacity for generating novel yet valid hypotheses under open-ended constraints. Methodologies such as Scideator Radensky et al. (2025) have been developed to investigate human-LLM collaboration to facilitate grounded idea and hypothesis generation. Other approaches have focused on ensuring the generated hypotheses are well-founded; for instance, HypER generates literature-grounded hypotheses with clear provenance Vasu et al. (2025), while O’Neill et al. (2025) leverage structured data from scientific papers for the same purpose. Within natural science, benchmarks Qi et al. (2024) and methods O’Brien et al. (2024) have extended hypothesis generation into the biomedical domain. Meanwhile, MOOSE-Chem Yang et al. (2025) offers a systematic evaluation benchmark and an agent framework specifically for hypothesis discovery in chemistry.
@@ 3.4 Data Analysis and Organization
In this stage, LLMs assist the scientific workflow by automating processes related to data organization, presentation, and analysis. For data presented in tabular format, Chain-of-Table Wang et al. (2024d) proposes a method to enhance tabular understanding by incorporating evolving tables within the reasoning chain of LLMs. Concurrently, Deng et al. (2024a) highlight the potential of integrating visual information to improve multimodal understanding, thereby aiding tabular comprehension. More recently, Wu et al. (2025) introduced TableBench, a comprehensive benchmark for table-based question answering under practical industrial scenarios.
@@
Beyond tabular data, charts represent another important format for organizing and storing information derived from experimental data. Early benchmarks, exemplified by ChartQA Masry et al. (2022), examined the capabilities of vision transformers in chart-based question answering. Subsequent works, including CharXiv Wang et al. (2024e) and ChartX Xia et al. (2025), have expanded the scope of chart understanding scenarios by utilizing human-curated chart generation or by incorporating real-world chart data sourced from arXiv preprints. Regarding chart generation, AutomaTikZ Belouadi et al. (2024) formulates the process as TikZ code generation from caption text and has demonstrated the efficacy of fine-tuning LLMs using open scientific figure data. More recently, Text2Chart31 Zadeh et al. (2025) employed reinforcement learning with automated feedback to refine chart generation capabilities within the Matplotlib library.
@@ 3.5 Conclusion and Hypothesis Validation
In this context, a significant focus of contemporary research involves investigating the utility of LLMs as reviewers for artificial intelligence papers. ReviewerGPT Liu and Shah (2023) initially explored the capability of LLMs to identify deliberately inserted errors within research papers, highlighting the necessity for more robust systems to conduct comprehensive reviews. Zhou et al. (2024a) further evaluated static LLMs in the context of reviewing real-world conference papers using a multiple-choice format. Du et al. (2024) conducted a comprehensive analysis of LLM review quality through extensive human studies and comparisons, revealing weaknesses in their ability to identify deficiencies. ClaimCheck Ou et al. (2025) further investigated the capabilities of LLMs in critiquing research claims, demonstrating that this task remains challenging even for highly advanced models such as OpenAI’s o1 OpenAI (2024). Beyond reviewing, other work has focused on the subsequent step of paper revision, with systems like XtraGPT enabling human-AI collaboration for controllable academic paper revisions Chen et al. (2025a). Concurrently, research highlights the potential to address these limitations by incorporating multi-agent systems with specialized roles Tyser et al. (2024); Xu et al. (2025a), through LLM alignment via reinforcement learning Weng et al. (2025), or by employing novel frameworks like generative adversarial reviews Bougie and Watanabe (2024).
@@
Another important application at this stage is the automatic validation of hypotheses by LLMs. Takagi et al. (2023) demonstrated that LLMs possess considerable capabilities in automatically generating code to verify research hypotheses within simplified machine learning problems. Benchmarks such as SciReplicate-Bench Xiang et al. (2025) and PaperBench Starace et al. (2025) have further extended this concept to evaluating the replication of real-world research papers. A distinct but related line of inquiry explores predicting empirical AI research outcomes directly with language models, assessing whether LLMs can anticipate experimental results without full execution Wen et al. (2025). Furthermore, Xu et al. (2025c) have navigated this domain into physics research, aiming to enhance the interpretability of the discovery process through the use of multi-agent workflows.
@@ 4.1 Machine Learning Research
Automated Machine Learning (AutoML) Shen et al. (2024) endeavors to generate high-performing modeling configurations for a given task in a data-driven manner. With the advent of LLM-based agents, several studies have explored their application in the automated modeling of machine learning (ML) tasks. A suite of benchmarks has emerged to track progress in this area. MLAgentBench Huang et al. (2024a) evaluates the capabilities of LLMs in designing and executing ML experiments, revealing that performance is often contingent upon task familiarity. Similarly, MLRC-Bench Zhang et al. (2025b) and RE-Bench Wijk et al. (2024) further probe the limits of these agents, assessing their ability to solve novel ML research challenges and comparing their R&D capabilities against human experts. MLGym Nathani et al. (2025) offers valuable resource and benchmark for advancing these AI research agents.
To address the challenges posed by these benchmarks, various agentic frameworks have been proposed. The IMPROVE framework Xue et al. (2025) highlighted the significance of iterative refinement mechanisms. CodeScientist Jansen et al. (2025) incorporated an ML modeling agent with machine-generated ideas, while BudgetMLAgent Gandhi et al. (2025) leveraged curated expert collaboration frameworks to achieve superior results with cost-effective models. More recent end-to-end systems like MLR-Copilot Li et al. (2024d) and the multi-agent framework MLZero Fang et al. (2025) aim for fully autonomous machine learning research and automation. Pushing the boundaries of automation even further, some work has explored the use of language models to directly propose LM architectures Cheng et al. (2025a), moving beyond orchestration to direct model creation.
@@ 4.2 Data Modeling and Analysis
Automated data-driven analysis, encompassing statistical data modeling and hypothesis validation, represents a foundational application area for LLM-assisted scientific discovery. InfiAgent-DABench Hu et al. (2024b) benchmarked the capabilities of LLMs in static code generation and execution for data analysis using CSV files. Subsequent benchmarks, such as BLADE Gu et al. (2024), DiscoveryBench Majumder et al. (2024), and DSBench Jing et al. (2024), have improved evaluation robustness by incorporating real-world research papers and expert-curated analytics to assess how far agents are from human expert performance. These studies indicate that most LLMs struggle with complex data analytics tasks, even when operating within an agent framework Zheng et al. (2023). To address these challenges, DS-Agent Guo et al. (2024b) proposes to enhance LLM performance by incorporating a case-based reasoning method to improve domain knowledge acquisition. In a related effort, DAgent Xu et al. (2025b) extended the application domain to querying relational databases and enabled report generation using results derived from decomposed sub-problems.
@@ 4.3 Function Discovery
Function discovery, which aims to identify the underlying equations from observational data of variables, has been significantly influenced by the advancement of AI-driven symbolic regression (SR) Udrescu and Tegmark (2020); Kamienny et al. (2022). To enhance this process, LLM-SR Shojaee et al. (2025a) leveraged the prior domain knowledge of LLMs and incorporated feedback from clustered memory storage, while DrSR Wang et al. (2025a) proposed a dual reasoning framework that utilizes both data and experience for scientific equation discovery. To systematically assess these capabilities, LLM-SRBench Shojaee et al. (2025b) introduced a benchmark for evaluating LLMs as function discovery agents, which incorporates function transformations to mitigate data contamination. Furthermore, other studies have explored the capabilities of LLMs in discovering complex models within specific domains, such as Physics Koblischke et al. (2025), Statistics Li et al. (2024b), and automated neural scaling law discovery Lin et al. (2025).
@@ 4.4 Natural Science Research
Research has largely focused on applying LLMs to autonomous research workflows for natural science discovery. Auto-Bench Chen et al. (2025b) evaluated LLMs on chemistry and social science tasks based on causal graph discovery, revealing that LLMs perform effectively only when task complexity is highly limited. In contrast, ScienceAgentBench Chen et al. (2025c) provided a multidisciplinary benchmark for LLMs operating within agent frameworks such as CodeAct Wang et al. (2024b) and self-debug Chen et al. (2023). This benchmark highlighted the necessity for tailored agent workflows for such explorative tasks.
@@ 4.5 General Research
Apart from specialized domain applications, some benchmarks have broadly evaluated diverse tasks from different stages of scientific discovery. DiscoveryWorld Jansen et al. (2024) created a virtual environment for LLM agents to conduct simplified scientific exploration. In Liu et al. (2025a), various application scenarios for AI agents in research were comprehensively discussed, supported by preliminary evaluation datasets. Similarly, CURIE Kon et al. (2025) proposed a benchmark and an agentic framework for rigorous and automated scientific experimentation. While EAIRA Cappello et al. (2025) focused on assessing the ability of LLMs to perform in a real-world research assistant role using various task formats.
@@ 6 Challenges and Future Directions
As LLM-based systems gain independent reasoning and action capabilities, their potential for risks—ranging from amplified societal biases to deliberate misuse like generating harmful substances or challenging human control—becomes increasingly salient and complex He et al. (2023); Ahadian and Guan (2024); Bengio et al. (2025). With AI capabilities and societal norms in constant flux, alignment is consequently an imperative, continuous process demanding adaptive governance and evolving value systems Li et al. (2024e). This requires embedding ethical constraints directly in scientific AI design frameworks, alongside vigilant oversight, to ensure advancements serve human well-being and the common good.
@@
Consequently, the scope of this survey has certain limitations. Firstly, we do not provide an exhaustive review of research focused on the development of general-purpose scientific LLMs for domain-specific reasoning or application. These areas, while crucial to the broader landscape of AI in science, are extensively covered in other existing surveys and fall outside our specific focus on the autonomy paradigm. Secondly, while we acknowledge the importance of fundamental LLM capabilities such as planning, code generation, and agentic decision-making, this survey does not delve deeply into orthogonal benchmarks or methodologies related to these general abilities. These exclusions were deliberate to maintain a focused exploration of the transformative roles and increasing independence of LLMs throughout the scientific research lifecycle.
@@
Acknowledgements
We thank all the anonymous reviewers and meta reviewers for their valuable comments. The authors of this paper were supported by the ITSP Platform Research Project (ITS/189/23FP) from ITC of Hong Kong, SAR, China, and the AoE (AoE/E-601/24-N), the RIF (R6021-20) and the GRF (16205322) from RGC of Hong Kong, SAR, China.